# VLM See, Robot Do: Human Demo Video to Robot Action Plan via Vision Language Model

### v1

@@ Abstract
Vision Language Models (VLMs) have recently been adopted in robotics for their capability in common sense reasoning and generalizability. Existing work has applied VLMs to generate task and motion planning from natural language instructions and simulate training data for robot learning. In this work, we explore using VLM to interpret human demonstration videos and generate robot task planning. Our method integrates keyframe selection, visual perception, and VLM reasoning into a pipeline. We named it SeeDo because it enables the VLM to “see” human demonstrations and explain the corresponding plans to the robot for it to “do”. To validate our approach, we collected a set of long-horizon human videos demonstrating pick-and-place tasks in three diverse categories and designed a set of metrics to comprehensively benchmark SeeDo against several baselines, including state-of-the-art video-input VLMs. The experiments demonstrate SeeDo’s superior performance. We further deployed the generated task plans in both a simulation environment and on a real robot arm. The code, demos, and data can be found at ai4ce.github.io/SeeDo.
@@ I Introduction
Large Vision Language Models (VLMs) or Multimodal Large Language Models (MLLMs) have been drawing significant interest in recent AI research. They have also been embraced by the robotics research community for their rich semantic knowledge and common-sense reasoning capabilities. Some research utilizes VLMs as an interface for parsing human language instructions to generate task plans [1, 2, 3, 4]. Some leverage VLMs to assist in motion and trajectory planning [5, 6, 7, 8], or incorporate VLMs in data generation systems to simulate real-world data for training robot policies [9, 10, 11]. These works typically use text, images, or both as inputs to the VLMs. In this work, we explore a different type of input modality: videos. We pose the question—can robots, with the help of VLMs, learn task plans directly from human demonstration videos?
@@
Figure 1: VLM See, Robot Do. We designed a pipeline centered around a large language vision model to interpret human demonstration videos and generate task plans for robots. The generated plans are converted to robotics code and deployed to simulated and real-world robots.
@@
While language instructions are efficient in many application scenarios, some tasks are inherently difficult to precisely express in plain language. Videos provide a more straightforward representation and are particularly suitable for long-horizon tasks with multiple steps or those that involve temporal or spatial dependencies. Moreover, videos are a natural learning medium for humans, who frequently acquire skills and parse task steps by observing demonstrations. Video data is abundant on the internet and holds the promise to be leveraged in scaling robot learning [12]. However, significant challenges remain in teaching robots to learn from human videos due to the substantial domain gap between robots and humans. While research in imitation learning tries to close the gap, long-horizon tasks often require collecting an increasing number of demonstrations. Motivated by this, we propose breaking down human demonstration videos with SeeDo. The SeeDo interprets a long-horizon human demonstration video into sub-task steps for robots, which can then be executed by language model programs (LMPs) [1] and low-level action primitives.
@@
Unlike directly feeding long-horizon human demonstration videos into imitation learning models, interpreting them with SeeDo offers several advantages due to the capabilities of VLMs. First, the rich commonsense knowledge of VLMs enables them to understand objects and their relations, allowing the model to grasp the task despite the embodiment gap. Second, their strong zero-shot generalization makes them more robust to environmental changes in the video. The plans generated by VLMs can remain valid even when the appearances or locations of objects, or the surroundings, differ between demonstration videos and real-world deployment.
@@
Despite the advantages of VLMs, we find that they struggle with processing every frame in a full-length video and accurately determining spatial relations, which are key features of videos in robotics applications. To address this, SeeDo is equipped not only with a VLM reasoning module but also with a keyframe selection module and a vision perception module, as shown in Figure 1. The keyframe selection identifies critical frames based on hand-speed heuristics, while the vision perception enhances the VLM’s ability to track objects, improving its overall perception. To test the full pipeline, we collected a benchmark of long-horizon pick-and-place tasks with human demonstration videos in three diverse categories: vegetable organization, garments organization, and wooden block stacking. All exhibit strong temporal and spatial dependencies. We evaluated SeeDo along with baselines such as the state-of-the-art video VLMs [13] and SeeDo gives the best performance.
In summary, the contributions of this work are:
@@
  • We propose SeeDo, a pipeline comprising keyframe selection, visual perception, and VLM reasoning modules, to interpret human demonstration videos and generate task plans for robots.
@@
  • We collected a benchmark of long-horizon human demonstration videos of pick-and-place tasks across three diverse categories and designed new evaluation metrics.
@@
  • We show that SeeDo achieves superior performance against the state-of-the-art baselines and the generated task plans can be successfully deployed to both simulation and real-robot scenarios.
@@ II Related Works
VLMs for task planning. Large language models have shown amazing emergent abilities and generalizability to new tasks since the release of GPT-3 [14], and have been used in generating task plans [15, 16, 17, 18]. To generate valid task planning, this body of work usually requires carefully curated instructions to prompt the VLMs [19] and often relies on access to advanced close-sourced VLMs such as GPT-4 [20] for good performance [18]. The task planning ability of VLMs has also been explored for robotics control [7, 15, 21, 5]. One line of work generates robot executable codes as the medium of task plans [1, 22]. In these works, the VLMs take human language instructions and sometimes images as inputs, reason the instructions, and output the task plans as function-calling codes, referred to as language model programs (LMPs), which call the wrapped API of robotics action primitives. [23] builds a video-language planning pipeline by using an embodied VLM [24] to break long-horizon language instructions into steps, prompt a video model to generate video rollouts of the future, and assess current progress. They all rely on human language instructions to specify the task while our SeeDo explores directly using real-world human demonstration videos as the task specification.
@@
VLMs for robot. With the help of strong common sense reasoning and rich semantic knowledge exhibited by the pretrained large VLMs [25], a line of literature [3, 26, 22] has demonstrated that robots can take non-expert human language instructions more effectively than before [27]. Robotics researchers also go beyond using pretrained VLMs and leverage robotics data to train vision-language-action models (VLAs) catered for direct robotics control [24, 28, 6]. By harnessing the rich common sense knowledge compressed in the large VLMs, useful data generation pipelines can be built to generate simulation data for training intelligent robots [9, 11]. In this work, we explore leveraging VLMs as a tutor to the robots, interpreting human videos into task plans that can be further executed via LMPs.fRobot learning from human videos. Demonstration videos provide a direct supervision signal for robot learning. Many existing works leverage teleoperated robotics videos to train robot policies via imitation learning [12, 29, 30]. For its massive amount compared to simulation and teleoperation, human video data has always held the promise of scaling up robot learning [31]. There has been a long interest in robot learning from human demonstrations from the early stage of robot learning [32, 33] to the recent deep imitation learning [34, 35, 36, 37]. Research in one-shot imitation learning [38, 39, 40, 41, 42] aims to learn robot policies from a single demonstration, but they are limited in generalizability and confined to short-horizon tasks. In general, having robots directly learn from human demonstration videos is still challenging due to the big domain gap between humans and robots. It also struggles with changes in the environment and object appearances between videos and deployments and often requires collecting many more demonstrations when the tasks become long-horizon [43]. To mitigate these challenges, [37] proposes to break down the imitation learning process into training a latent planner to predict hand trajectory from human play data and training the planner-guided imitation policies on robotics data. While sharing a similar motivation of decomposing planning and action learning, this work focuses on the planning part and takes a different approach. SeeDo leverages pretrained VLMs to directly interpret human demonstration videos into textual plans and the generated plans are processed into LMPs to call any action primitives whether they are trained-based, control-based, or pre-programmed.
@@
VLMs for video input. Recent VLMs have been trained to accommodate multiple modalities of inputs including videos [44, 13, 45, 46, 47, 48] and can do video analysis tasks like question answering (QA) and video captioning. Video analysis [49] has also been included as part of the benchmark set to evaluate VLMs. Our approach resembles a video QA or captioning setup but is grounded specifically in robotics scenarios. We also included several top-performing commercial and open-source VLMs on the VideoMME [49] benchmark as our baselines.
@@ III Method
Figure 2: The SeeDo pipeline consists of three modules. From left to right, a) The Keyframe Selection module detects the operating hand in the video and plots its speed. The speed valleys are identified as keyframes. b) The Visual Perception module detects and tracks objects and then applies the tracking results as visual prompts to each keyframe. c) The VLM Reasoning module instructs the GPT-4o to interpret keyframes, identify objects and actions in each keyframe, and generate task plans from the demonstration video. d) Plan execution. The generated task plans are processed by Code-as-Policies into language model programs (LMPs) and call the robot APIs for execution.
@@
In preliminary studies, we find that simply instructing plain VLMs with human demonstration videos yields poor results. They constantly struggle with processing and retaining information from all frames, often confusing the temporal order and spatial relationships of objects. Therefore, inspired by [4] we took a system-first approach and designed a pipeline centered around the VLM to enhance its capabilities. As a result, SeeDo comprises three modules: Keyframe Selection module, Visual Perception module, and VLM Reasoning module. The full pipeline is illustrated in Figure 2 and we detail each module below.
@@
Keyframe Selection. Context length is a major constraint when VLMs process videos. Open-source VLMs often simply sample frames uniformly [13, 45], but this approach is less effective for demonstration videos, as frames showing important actions may be missed. Therefore, we adopt a heuristic approach that selects keyframes by detecting hand speed. Hand-object interactions are critical in demonstration videos [34, 35] and we observe that hands typically move slower when picking or placing objects, providing a clue to locating keyframes. Similar observations have also been made in the literature [50]. Specifically, we leverage a lightweight method [51] to detect hand and plot hand speed over time [52]. The resulting plot is interpolated and smoothed, producing a wave-like representation of hand speed, with the frames corresponding to the troughs selected as keyframes. Since hand detection is not always perfect and some troughs may be a result of interpolation, we further filter out noise keyframes. In the VLM reasoning module, we prompt the VLM to assess if the frames contain hand-object interactions, which yields reasonably accurate results.
@@
Visual Perception. Visual shortcomings of current VLMs have been reported in the literature [53]. We also find that VLMs often struggle to accurately determine object locations and their spatial relations. They also frequently fail to consistently differentiate between visually similar objects over time, which are crucial for planning tasks from human demonstration videos. To address these issues, we introduce a visual perception module in SeeDo to enhance the VLM’s visual capabilities. Specifically, we first instruct the VLM to identify objects in the video, and then use an open-vocabulary object detector [54] to extract object bounding boxes in the first frame. These bounding boxes serve as prompts to the most recent Segment Anything Model (SAM2) [55] for video tracking. The resulting tracking IDs and mask contours are annotated onto the previously selected keyframes as visual prompts [56]. The prompted keyframes are subsequently given to the VLM Reasoning module.
@@
Figure 3: We collected human demonstration videos across three diverse categories as our tasks and carried out both simulation and real-world experiments. Tasks from left to right: vegetable organization, garment organization, and wooden block stacking.
@@
VLM Reasoning. The VLM Reasoning module uses chain-of-thought (CoT) prompting [19] to generate task planning steps as the final output of SeeDo. For the specific model selection, we choose GPT-4o as the VLM in SeeDo for its superior performance. We note that using commercial, closed-source models is not uncommon in the research community. Previous work in VLM for robotics [3, 2, 1, 57, 21, 58] has mostly adopted the GPT series or other close-sourced models such as the PALM [59, 24]. A key design in this module is the incorporation of visual prompts in the reasoning. Previous studies have shown that providing VLMs with images containing prompted visual cues is effective for manipulation [60] and navigation tasks [61]. In SeeDo, identifying objects and understanding their spatial relationships is crucial for interpreting human demonstration videos, and this is greatly enhanced by the visual prompts derived from video tracking in the Visual Perception module. Specifically, we maintain an object list obtained from the Visual Perception module in the prompt. The mask contours and tracking IDs are used as visual prompts in the keyframes to aid object identification. By using contours instead of full masks we ensure that the objects of interest are highlighted without obstructing their appearance. The center coordinates of the masks, combined with the corresponding tracking IDs, are then appended to the textual prompts to imply the VLM of the objects’ spatial relations. Empirically, we find this design particularly helpful when the human demonstration videos involve visually similar objects, such as when playing with uncolored wooden blocks. Below we provide an example CoT prompt in SeeDo. The complete prompt will be made available along with the code.
@@
Plan execution. The SeeDo-generated task plans can be seamlessly processed step by step by any robot action model that can take text input. Specifically, following the approaches in [1] and [21], we use Language Model Programs (LMPs) to implement the task plans on a UR10e robot arm in both Pybullet simulations [62] and real-world deployment.
@@ IV Experiments
LLaVA [13] 0.00 0.00 2.75 0.00 0.00 31.61 0.00 0.00 2.56
VILA [45] 5.26 5.26 14.38 0.00 0.00 0.90 0.00 0.00 3.41
@@
We collected human demonstration videos across three diverse long-horizon pick-and-place tasks and designed a set of metrics to evaluate the success rate. We first compare SeeDo to baselines across all three tasks. Then, we present ablation studies to assess the impact of the module design. Finally, we analyzed and discussed the types of errors that occurred with SeeDo and the baselines. The experiments were conducted on a fixed robotic arm platform using a UR10E robot arm. Qualitative results are shown in Figure 4.

### v2

@@ Abstract
Vision Language Models (VLMs) have recently been adopted in robotics for their capability in common sense reasoning and generalizability. Existing work has applied VLMs to generate task and motion planning from natural language instructions and simulate training data for robot learning. In this work, we explore using VLM to interpret human demonstration videos and generate robot task planning. Our method integrates keyframe selection, visual perception, and VLM reasoning into a pipeline. We named it SeeDo because it enables the VLM to “see” human demonstrations and explain the corresponding plans to the robot for it to “do”. To validate our approach, we collected a set of long-horizon human videos demonstrating pick-and-place tasks in three diverse categories and designed a set of metrics to comprehensively benchmark SeeDo against several baselines, including state-of-the-art video-input VLMs. The experiments demonstrate SeeDo’s superior performance. We further deployed the generated task plans in both a simulation environment and on a real robot arm. The code, demos, prompts and data can be found at ai4ce.github.io/SeeDo.
@@ I Introduction
Large Vision Language Models (VLMs) have drawn significant interest in recent robotics and embodied AI research for their rich semantic and common-sense knowledge. Some research utilizes VLMs as an interface for parsing human language instructions to generate task plans [1, 2, 3, 4]. Some employ pretrained VLMs for further fine-tuning to learn the mapping from visual inputs and language instructions to actions [5, 6], or leverage the general knowledge of VLMs to identify salient objects or key points [7, 8].
@@
Figure 1: VLM See, Robot Do. We designed an agent framework centered around a large Vision Language Model to interpret long-horizon human demonstration videos into task plans in natural language, which are then executed in simulated and real-world robots via language model programs and action primitive functions.
@@
Although language instructions are effective in many scenarios, some tasks are inherently difficult to describe precisely using plain language. Additionally, as tasks become longer-horizon, providing instructions in language can become tedious and inefficient. In contrast, videos offer a more straightforward medium, making them particularly well-suited for long-horizon tasks that involve multiple steps or require understanding of temporal and spatial dependencies. Moreover, video data is abundant on the internet and holds the promise to be leveraged in scaling robot learning [9]. However, significant challenges remain in teaching robots to learn from human videos due to the substantial domain gap between robots and humans. While research in imitation learning tries to close the gap, long-horizon tasks often require collecting an increasing number of demonstrations. In contrast, humans can effectively learn new tasks by watching a few demonstration videos and intuitively breaking them down into sequential steps. Inspired by this capability, we propose SeeDo, a modularized agent centered around a VLM. SeeDo is designed to interpret long-horizon human demonstration videos into sub-task steps in natural language, which can then be executed by language model programs (LMPs) [1] and low-level action primitives.
@@
Compared to directly feeding long-horizon human demonstration videos into imitation learning models, interpreting them with SeeDo offers several advantages thanks to the capabilities of VLMs. First, VLMs’ rich commonsense knowledge enables them to understand objects and their relationships, helping robots understand the task goals despite the embodiment gap. Second, their strong zero-shot generalization ability makes them more robust to unseen environments and objects during execution. The plans generated by VLMs remain valid even when object appearances, locations, or the surroundings differ between demonstration videos and real-world deployment.
@@
Despite these advantages, we find VLMs struggle with processing a large number of video frames. They also lack the ability to accurately recognize objects and their spatial relations from the video, which is key information in robotics applications. To mitigate these limitations, SeeDo integrates not only with a VLM interpreter module but also with a keyframe selection module and a visual prompting module, as illustrated in Figure 1. The keyframe selection module identifies critical frames based on hand-speed heuristics, while the visual prompting module enhances the VLM’s ability in visual and spatial perception. To evaluate SeeDo, we curated a benchmark of long-horizon pick-and-place tasks with human demonstration videos in three diverse categories: vegetable organization, garments organization, and wooden block stacking. All exhibit strong temporal and spatial dependencies. We compare SeeDo against baselines such as the state-of-the-art video VLMs [10, 11], and find that SeeDo achieve the best performance.
In summary, the contributions of this work are as follows:
@@
  • We introduce SeeDo, a VLM-based agent that integrates keyframe selection, visual prompting, and VLM interpreter modules to interpret long-horizon human demonstration videos and generate task plans for robots.
@@
  • We curate a benchmark of long-horizon human demo videos of pick-and-place tasks across three diverse categories and propose new evaluation metrics.
@@
  • We show that SeeDo achieves superior results compared to state-of-the-art baselines that directly process videos or images as input. Furthermore, the generated task plans can be successfully deployed in both simulation and real-robot scenarios.
@@ II Related Works
VLMs for task planning. Large language models have shown amazing emergent abilities and generalizability to new tasks since the release of GPT-3 [12], and have been used in generating task plans [13, 14, 15, 16]. To generate valid task planning, this body of work usually requires carefully curated instructions to prompt the VLMs [17] and often relies on access to advanced close-sourced VLMs such as GPT-4 [18] for good performance [16]. The task planning ability of VLMs has also been explored for robotics control [19, 13, 20]. One line of work generates robot executable codes as the medium of task plans [1, 21]. In these works, the VLMs take human language instructions and sometimes images as inputs, reason the instructions, and output the task plans as function-calling codes referred to as language model programs (LMPs), which call the wrapped API of robotics action primitives. [22] builds a video-language planning pipeline by using an embodied VLM [23] to break long-horizon language instructions into steps, prompt a video model to generate video rollouts of the future, and assess current progress. They rely on human language instructions to specify the task while SeeDo explores directly interpreting long-horizon real-world human demonstration videos into the task specification.
VLMs for robots. With the help of strong common sense understanding and rich semantic knowledge exhibited by the pretrained VLMs [24], a line of literature [3, 25, 21] has demonstrated that robots can take non-expert human language instructions more effectively than before [26]. Robotics researchers also go beyond using pretrained VLMs and leverage robotics data to train vision-language-action models (VLAs) catered for direct robotics control [23, 27, 5]. Harnessing the rich common sense knowledge compressed in the large VLMs, useful data generation pipelines can be built to generate simulation data for training intelligent robots [28, 29]. By leveraging the semantic knowledge of VLM, important objects or key points for execution can be selected and assigned affordance or constraint functions, thereby transforming the robot execution problem into an optimization problem over these functions. [7, 8]. In this work, we explore leveraging VLMs as an agent to the robots, interpreting human videos into task plans that can be further executed via LMPs.
@@
Robot learning from human videos. Demonstration videos provide a direct supervision signal for robot learning. Many existing works leverage teleoperated robotics videos to train robot policies via imitation learning [9, 30, 31]. For its massive amount compared to simulation and teleoperation, human video data has always held the promise of scaling up robot learning [32]. There has been a long interest in robot learning from human demonstrations from the early stage of robot learning [33, 34] to the recent deep imitation learning [35, 36, 37, 38]. Research in one-shot imitation learning [39, 40, 41, 42, 43, 44] aims to learn robot policies from a single demonstration, but they are limited in generalizability and confined to short-horizon tasks. In general, having robots directly learn from human demonstration videos is still challenging due to the big domain gap between humans and robots. It also struggles with changes in the environment and object appearances between videos and deployments and often requires collecting many more demonstrations when the tasks become long-horizon [45]. To mitigate these challenges, [38] proposes to break down the imitation learning process into training a latent planner to predict hand trajectory from human play data and training the planner-guided imitation policies on robotics data. While sharing a similar motivation of decomposing planning and action learning, our work focuses on the planning part and takes a different approach. SeeDo leverages pretrained VLMs to directly interpret human demonstration videos into text plans and the generated plans are processed into LMPs to call any action primitives whether they are trained-based, control-based, or pre-programmed.
@@
VLMs for video understanding. Recent VLMs have been trained to accommodate multiple modalities of inputs including videos [11, 46, 47, 10] and can do video analysis tasks like question answering (QA) and video captioning. Video analysis [48] has also been included as part of the benchmark set to evaluate VLMs. Our approach resembles a video QA or captioning setup but is grounded specifically in robotics scenarios. We also included several top-performing proprietary and open-source VLMs on the VideoMME [48] benchmark as our baselines.
@@ III Method
Figure 2: The SeeDo agent consists of three modules. From left to right, a) The Keyframe Selection module detects the operating hand in the video and plots its speed. The speed valleys are identified as keyframes. b) The Visual Prompting module detects and tracks objects and then applies the tracking results as visual prompts to each keyframe. c) The VLM Interpreter module instructs the GPT-4o to interpret keyframes, identify objects and actions in each keyframe, and generate task plans from the demonstration video. d) Plan Execution. The generated task plans are processed by Code-as-Policies into language model programs (LMPs) and call the robot APIs for execution.
@@
In preliminary studies, we find that simply instructing pretrained VLMs with human demonstration videos yields poor results. They constantly struggle with processing and retaining information from all frames, often confusing objects’ temporal order and spatial relationships. Therefore, we design an agent framework centered around the VLM to enhance its capabilities. As a result, SeeDo comprises three modules: Keyframe Selection module, Visual Prompting module, and VLM Interpreter module. The full pipeline is illustrated in Figure 2 and we detail each module below.
@@
Keyframe Selection. Context length becomes a major constraint when VLMs process long-horizon videos. Open-source VLMs often simply sample frames uniformly [10, 46], but this approach is less effective for long-horizon demonstration videos, as frames showing important actions may be missed. To overcome the context limit yet still capture critical information, we adopt a heuristic approach that selects keyframes by detecting hand speed. Hand-object interactions are critical in demonstration videos [35, 36] and we observe that hands typically move slower when picking or placing objects, providing a clue to locating keyframes [49]. Specifically, we use a lightweight model [50] to detect hand keypoints, and plot the speed of the keypoints’ center over time following [51]. The resulting plot is linearly interpolated and smoothed, producing a wave-like plot of hand speed. The frames corresponding to the troughs of the wave selected as keyframes. Since hand detection is not always perfect and some troughs may be a result of interpolation, we further filter out irrelevant keyframes. In the VLM planner module, we prompt the VLM to filter out the frames showing no hand-object interactions, which yields reasonably accurate results.
@@
Visual Prompting. Visual shortcomings of current VLMs have been reported in the literature [52]. We also find that VLMs often struggle to accurately determine object locations and their spatial relationships. They also frequently fail to consistently differentiate between visually similar objects over time, which is crucial for task planning from long-horizon human demonstration videos. To alleviate these issues, we introduce a visual prompting module in SeeDo that enhances the visual capabilities of the VLM. The module first instructs the VLM to identify objects in the frames and then use an open-vocabulary object detector [53] to extract object bounding boxes in the first frame. These bounding boxes are prompted to the Segment Anything Model (SAM2) [54] for video tracking. The resulting tracking IDs and mask contours are rendered to the previously selected keyframes as visual prompts [55]. These visual-prompted keyframes are subsequently given to the VLM Interpreter module.
@@
Figure 3: We collect long-horizon human demonstration videos across three diverse categories as our benchmark and carry out both simulation and real-world experiments. Tasks from left to right: vegetable organization, garment organization, and wooden block stacking.
@@
VLM Interpreter. The VLM Interpreter module leverages chain-of-thought (CoT) [17] to generate task planning steps for robot execution. Any VLM can fit into the module, we choose GPT-4o for its popularity and widely recognized performance. We note that using closed-source models is not uncommon in the research community. Previous work in VLM for robotics [3, 2, 1, 56, 20, 57] has mostly adopted the GPT series or other close-sourced models such as the PALM [58, 23]. A key design is the incorporation of visual prompts. Previous studies have shown that providing VLMs with images containing prompted visual cues is effective for manipulation [59] and navigation tasks [60]. In SeeDo, identifying objects and understanding their spatial relationships is crucial to interpret long-horizon human demonstration videos, and this is greatly enhanced by the visual prompts derived from video tracking in the Visual Prompting module. Specifically, we maintain an object list obtained from the Visual Prompting module in the text prompt. The mask contours and tracking IDs are used as visual prompts in the keyframes to aid object identification. By using contours instead of full masks we ensure the objects of interest are highlighted without obstructing their appearance. The center coordinates of the masks, combined with the corresponding tracking IDs, are then appended to the textual prompts to imply the VLM of the objects’ spatial relationships. We find this design particularly helpful when the human demonstration videos involve visually similar objects, such as when playing with uncolored wooden blocks. Below is an example CoT prompt in SeeDo. The complete prompt will be made available along with the code.
@@
Plan Execution. The SeeDo-generated task plans can be seamlessly processed step by step by any robot action model that can take text input. Specifically, following the approaches in [1] and [20], we use Language Model Programs (LMPs) to implement the task plans on a UR10e robot arm in both Pybullet simulations [61] and real-world deployment. Action primitives are predefined as functions and can be called by the LMPs. In real-world experiment, we follow [1, 20] and first use a segmentation model to segment all objects of interest in the RGB images, then query the depth image with the same coordinates to acquire 3D information of the objects.
@@ IV Experiments
LLaVA-OneVision [10] 0.00 0.00 2.75 0.00 0.00 31.61 0.00 0.00 2.56
LLaVA-NeXT-Video-7B [47] 0.00 0.00 9.03 0.00 0.00 4.00 0.00 0.00 3.85
VILA1.5-8B [46] 5.26 5.26 14.38 0.00 0.00 0.90 0.00 0.00 3.41
@@
We curated long-horizon human demonstration videos across three diverse pick-and-place tasks and designed a set of metrics for comprehensive evaluation. In addition to comparing SeeDo to baselines across all three tasks. We also present ablation studies to assess the impact of separate modules. Finally, we analyzed and discussed the types of errors that occurred with SeeDo and the baselines. Qualitative results are shown in Figure 4.